Diffusion Morning Report 01/10/06

Ted Berg tberg at snf.stanford.edu
Tue Jan 10 06:24:19 PST 2006


Good Morning All,
    Looking good this morning.  All of the LPs except TEOS2 are back up 
and diffusion is looking good.  TEOS2 tube was pulled and cleaned
while O2 MFC is being repaired.  Everything else looks OK for now.

  	Date: 	1/10/2006 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	13.5/30
Tylan 9-Poly 	Up 	  	9.77/30
Tylan 10-Nitride 	Up 	  	16.89/30 
Tylan 11-BPSG-LTO 	Up
	
	18.33/30
Tylan 12-FGA 	Up 	  	 
  	  	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	Up 	  	 
Thermco Nitride 	Down 	Testing 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	  
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	A on line 460 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20060110/c8871ddd/attachment.html>


More information about the diffusion mailing list